OpenStudy (anonymous):

how about it... f(x)=x^2+x and g(x)=2/(x+3) find value from : (f/g) (1)

OpenStudy (anonymous):

\[\Big(\frac{f}{g}\Big)(1) = \frac{f(1)}{g(1)}\]\[f(1) = (1)^2+(1) = 1+1 = 2\]\[g(1) = \frac{2}{(1)+3} = \frac{2}{4} = \frac{1}{2}\] So: \[\frac{f(1)}{g(1)} = \frac{2}{\frac{1}{2}} = 2*\frac{2}{1} = 2*2 = 4\]